Hello I have this word problem and i'm really confused on how to to write the equation for it. Here's the problem(Sorry for its length): A restaurant manager bought 20 packages of bagels. Some packages contained 6 bagels each, and the rest contained 12 bagels each. There were 168 bagels in all.How many packages of 12 bagels did the manager buy?
Answer by fcabanski(1391)

Let s= number of six bagel packages. Let t=number of twelve package bagels.


s + t = 20 since there are a total of 20 packages, which is the sum of the number of different types of packages.


6s + 12t = 168, Each six bagel package contributes 6 bagels for a total of 6s bagels. Each twelve bagel package contributes 12 bagels for a total of 12t bagels. The sum is the total number of bagels.


Solve the first equation for s in terms of t. s = 20 - t.


Substitute that into the second equation, and solve for t.


6(20-t) + 12t = 168


120 - 6t + 12t = 168


6t = 48 (And Mr. T = tough, fool).


t = 8


The manager bought 8 twelve bagel packages.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)